# Role of internal space correlations in the dynamics of a higher-dimensional Bianchi type-I universe: shear scalar and Hubble parameter perspectives

### v1

@@ I Introduction
Within four-dimensional general relativity (GR), expansion anisotropy—quantified by shear scalar ($\sigma^{2}$)—resembles the kinetic energy contribution of a canonical scalar field having an EoS parameter $w=1$—causality limit [7] for any physical source—thereby expansion anisotropy is the fastest diluting component with almost $\rho_{\sigma^{2}}\propto a^{-6}$ compared to any material sources having $w<1$ as the Universe expands, therefore we are not expecting to see the effects of anisotropy on the late universe atop of the $\Lambda$CDM (Lambda-cold dark matter) model. Generalized Friedmann equation containing the average Hubble parameter along with the shear scalar term retaining isotropic spatial curvature has been studied  [9, 10, 8, 7, 11, 12] and is well constrained [11, 12] in which a very small amount of present-day expansion anisotropy can not be excluded such that the data favors the isotropic expansion, with $\Omega_{\sigma 0}\leq 10^{-18}$ at $95\%$ confidence level from cosmic microwave background (CMB) data with lensing. However, an observational deviation from its stiff-fluid character in GR may imply a necessity for replacing GR by modified theories of gravity [^1] Modified gravity theories can be described as an imperfect fluid exerting different pressures in three directions (corresponding to different shear scalar evolutions) in this approach where the field equations are considered as effective Einstein equations, may shed light on dark energy, [16, 17, 18, 19, 20].. Similarly, the presence of extra dimensions in GR and in its alternatives requiring higher dimensions provide a fluid of geometric DE whose density parameter is determined by the number of extra dimensions, expansion anisotropy, if exists and can be detected, provides an alternative way of probing DE. For instance, the discovery of the expansion anisotropy in a very particular form different than GR predictions would be a strong signal in favor of modified theories and/or higher dimensions rather than the presence of a DE (such as $\Lambda$ and scalar fields) ingredient of the Universe assuming that GR is valid, see  [13]. For instance, in  [14], the higher dimensional Robertson-Walker cosmologies has been studied and then in a similar manner, its anisotropic extension [15] reveals that the shear scalar of the external space mimics a negative cosmological constant instead of stiff fluid-like evolution in the standard $\Lambda$CDM model under a higher dimensional steady-state universe ansatz. Interestingly, a higher dimensional steady-state universe gives rise to mathematically exactly the same Friedmann equation of the standard $\Lambda$CDM model for the external space, in which the higher-dimensional negative cosmological constant plays the role of the four-dimensional positive cosmological constant. In this study, we show that by allowing a negative correlation between the spaces different than [5], we are able to reproduce a higher dimensional steady-state universe via the kinematical constraint as well. For instance, in [13] it was shown that if one stays loyal to zero inertial mass density $p_{\Lambda}+\rho_{\Lambda}=0$ for DE on average allowing expansion anisotropy, shear scalar tracks the DE deforming it meaning that all cosmologies based on the canonical scalar fields should be reconsidered from this point of view, on top of that, along with a bonus, the kinetic term of the scalar field is replaced by an real observable, the shear scalar.
@@ III.1 Positive correlation of spaces ($\lambda>0$)
where $a_{1}$, $b_{1}$, $c_{1}$, $k_{1}$ and $k_{2}$ are positive integration constants satisfying $v_{\rm ext1}=(a_{1}b_{1}c_{1})^{1/3}$. Accordingly, the scale factor of internal space reads
@@ V Conclusion
Another interesting feature is that if the spatial coordinate changes signature; namely the space coordinate (of internal/external spaces) turns into time coordinate, scale factors in front ($a^{2}$ or $s^{2}$) remains invariant. Therefore, we could not distinguish this signature change on spatial coordinates in Einstein field equations for flat ($k=0$) space-like sections. Because Hubble parameter and its derivatives—as these contribute with the ratio form to field equations—remains the same leaving Einstein field equations exactly the same, on the other hand, curvature of space, if exists, contributes to field equations with ±$\pm 1/a^{2}$ (or ±$\pm 1/s^{2}$) depending on being closed/open spaces. A unique indication of the signature change on spatial coordinates is the spatial curvature, $k_{\rm ext}/a^{2}$ or $k_{\rm int}/s^{2}$. In [74], classical imaginary space extensions of Lorentzian theory with $a^{2}<0$ has been studied and cosmological constant switches its sign via a metric signature change across boundaries with a degenerate metric in Einstein-Cartan and Plebanski formulations of general relativity. Similarly in higher dimensional models, we realized that with a sudden change of the scale of the static internal space in the presence of internal curvature ($k_{\rm int}$) is realized as a sudden change of the cosmological constant $\tilde{\Lambda}_{\rm eff}=\tilde{\Lambda}-\frac{n(n-1)k_{\rm int}}{s^{2}}$—with accompanying abrupt change on the gravitational coupling constant—as observed in the external manifold, a detailed investigation is in progress and will be presented in a future work under progress. A flow scheme of the study is given in Fig. 1 in Appendix A to assist readers in navigating the paper and in relating with the other our studies in literature.

### v2

@@ I Introduction
Within four-dimensional general relativity (GR), expansion anisotropy—quantified by shear scalar ($\sigma^{2}$)—resembles the kinetic energy contribution of a canonical scalar field having an EoS parameter $w=1$—causality limit [7] for any physical source—thereby expansion anisotropy is the fastest diluting component with almost $\rho_{\sigma^{2}}\propto a^{-6}$ compared to any material sources having $w<1$ as the Universe expands, therefore we are not expecting to see the effects of anisotropy on the late universe atop of the $\Lambda$CDM (Lambda-cold dark matter) model. Generalized Friedmann equation containing the average Hubble parameter along with the shear scalar term retaining isotropic spatial curvature has been studied  [9, 10, 8, 7, 11, 12] and is well constrained [11, 12] in which a very small amount of present-day expansion anisotropy can not be excluded such that the data favors the isotropic expansion, with $\Omega_{\sigma 0}\leq 10^{-18}$ at $95\%$ confidence level from cosmic microwave background (CMB) data with lensing. As stated in [13], the anisotropic generalization of the $\Lambda$CDM model, under the stringent constraints, is not sufficient to solve the Hubble tension as well. However, an observational deviation from its stiff-fluid character in GR may imply a necessity for replacing GR by modified theories of gravity. [^1] Modified gravity theories can be described as an imperfect fluid exerting different pressures in three directions (corresponding to different shear scalar evolutions) in this approach where the field equations are considered as effective Einstein equations, may shed light on dark energy, [17, 18, 19, 20, 21].. Similarly, the presence of extra dimensions in GR and in its alternatives requiring higher dimensions provide a fluid of geometric DE whose density parameter is determined by the number of extra dimensions, expansion anisotropy, if exists and can be detected, provides an alternative way of probing DE. For instance, the discovery of the expansion anisotropy in a very particular form different than GR predictions would be a strong signal in favor of modified theories and/or higher dimensions rather than the presence of a DE (such as $\Lambda$ and scalar fields) ingredient of the Universe assuming that GR is valid, see  [14]. For instance, in  [15], the higher dimensional Robertson-Walker cosmologies has been studied and then in a similar manner, its anisotropic extension [16] reveals that the shear scalar of the external space mimics a negative cosmological constant instead of stiff fluid-like evolution in the standard $\Lambda$CDM model under a higher dimensional steady-state universe ansatz. Interestingly, a higher dimensional steady-state universe gives rise to mathematically exactly the same Friedmann equation of the standard $\Lambda$CDM model for the external space, in which the higher-dimensional negative cosmological constant plays the role of the four-dimensional positive cosmological constant. In this study, we show that by allowing a negative correlation between the spaces different than [5], we are able to reproduce a higher dimensional steady-state universe via the kinematical constraint as well. For instance, in [14] it was shown that if one stays loyal to zero inertial mass density $p_{\Lambda}+\rho_{\Lambda}=0$ for DE on average allowing expansion anisotropy, shear scalar tracks the DE deforming it meaning that all cosmologies based on the canonical scalar fields should be reconsidered from this point of view, on top of that, along with a bonus, the kinetic term of the scalar field is replaced by an real observable, the shear scalar.
@@ III.1 Positive correlation of spaces ($\lambda>0$)
where $a_{1}$, $b_{1}$, $c_{1}$, $k_{1}$ and $k_{2}$ are positive integration constants satisfying $v_{\rm ext1}=(a_{1}b_{1}c_{1})^{1/3}$. This solution has the same mathematical form with the four-dimensional anisotropic generalization of the de Sitter Universe, studied in [58] and was obtained from the asymp- totic form of the Kasner solution for large $t$ values in the presence of the cosmological constant.
Accordingly, the scale factor of internal space reads
@@ V Conclusion
Another interesting feature is that if the spatial coordinate changes signature; namely the space coordinate (of internal/external spaces) turns into time coordinate, scale factors in front ($a^{2}$ or $s^{2}$) remains invariant. Therefore, we could not distinguish this signature change on spatial coordinates in Einstein field equations for flat ($k=0$) space-like sections. Because Hubble parameter and its derivatives—as these contribute with the ratio form to field equations—remains the same leaving Einstein field equations exactly the same, on the other hand, curvature of space, if exists, contributes to field equations with ±$\pm 1/a^{2}$ (or ±$\pm 1/s^{2}$) depending on being closed/open spaces. A unique indication of the signature change on spatial coordinates is the spatial curvature, $k_{\rm ext}/a^{2}$ or $k_{\rm int}/s^{2}$. In [76], classical imaginary space extensions of Lorentzian theory with $a^{2}<0$ has been studied and cosmological constant switches its sign via a metric signature change across boundaries with a degenerate metric in Einstein-Cartan and Plebanski formulations of general relativity. Similarly in higher dimensional models, we realized that with a sudden change of the scale of the static internal space in the presence of internal curvature ($k_{\rm int}$) is realized as a sudden change of the cosmological constant $\tilde{\Lambda}_{\rm eff}=\tilde{\Lambda}-\frac{n(n-1)k_{\rm int}}{s^{2}}$—with accompanying abrupt change on the gravitational coupling constant—as observed in the external manifold, a detailed investigation is in progress and will be presented in a future work under progress [77]. A flow scheme of the study is given in Fig. 1 in Appendix A to assist readers in navigating the paper and in relating with the other our studies in literature.